# Counting Distinct Elements in the Turnstile Model with Differential Privacy under Continual Observation

### v1

@@ 1.2 Our techniques
We describe our algorithmic ideas via the reduction from distinct elements to the summation problem used in previous works [9, 30]. A mechanism for the summation problem outputs at every time step $t\in[T]$ the sum of the first $t$ elements of the stream. Dwork et al [23] and Chan et al. [13] use the binary tree mechanism to obtain a $O(\operatorname{polylog}T)$-accurate mechanism for summation. Given an input stream $x$ of length $T$ (to the $\mathsf{CountDistinct}$ problem), define a corresponding summation stream $s_{x}\in\{-1,0,1\}^{T}$. At time step $t\in[T]$, the entry $s_{x}[t]$ equals the difference in the count of distinct elements between time steps $t-1$ and $t$, i.e., $s_{x}[t]=\mathsf{CountDistinct}(x)[t]-\mathsf{CountDistinct}(x)[t-1]$. Then $\mathsf{CountDistinct}(x)[t]$ is precisely the sum of the first $t$ elements of $s_{x}$. In the insertion-only model, changing one entry of $x$ changes at most $2$ entries of $s_{x}$, and thus using the binary-tree mechanism together with a group privacy argument gives a mechanism with $O(\operatorname{polylog}T)$ additive error. For turnstile streams, even under the weaker notion of event-level privacy, a change in the stream $x$ can cause $\Omega(T)$ changes to $s_{x}$. To see this, consider the stream consisting of consecutive insertions ($+u$) and deletions ($-u$) of a single element $u\in\mathcal{U}$, and its event-neighboring stream where the first occurrence of $+u$ is replaced with $\perp$. This example illustrates that the difficulty of the distinct elements problem for turnstile streams lies with items that switch from being present to absent multiple times over the course of the stream. We present a mechanism that outputs private estimates of the number of distinct elements in a turnstile stream with optimal accuracy in terms of maximum flippancy.
@@
Our first key idea allows us to obtain a mechanism (Algorithm 1) that is given as input a flippancy upper bound $w$. For streams whose maximum flippancy is bounded by $w$, a change in $x$ causes at most $2w$ changes to $s_{x}$ (this is true for both item- and event-neighboring streams). This observation, combined with a group privacy argument, gives a mechanism with error $O(w\cdot\operatorname{polylog}\ T)$ directly from the error of the binary tree mechanism for summation. Previous works in the insertion-only setting [9, 30] use precisely this approach, setting $w=1$. To obtain the better $\sqrt{w}$ dependence on $w$ in our upper bound, we “open up” the analysis of the binary tree mechanism. By examining the information stored in each node of the binary tree for the summation stream, we show that changing the occurrences of one item in a stream $x$ with maximum flippancy $w$ can change the values of at most $w$ nodes in each level of the binary tree. The $\sqrt{w}$ dependence in the error then follows from the privacy guarantees of the Gaussian mechanism for approximate differential privacy. While our mechanism is only accurate for streams with maximum flippancy $w$, it must be private even for streams that violate this condition. To achieve this, our mechanism pre-processes the stream so that items are ignored after their flippancy exceeds $w$. This pre-processing can be done online.
@@
To avoid the need for an a-priori bound on $w$, we design a mechanism (Algorithm 3) that (privately, approximately) keeps track of the maximum flippancy of the prefix of the stream seen so far and invokes our first mechanism (Algorithm 1) with the current estimated maximum flippancy as an input. We track whether the maximum flippancy has doubled via the sparse vector algorithm [25]. The sparse vector algorithm receives queries of low constant sensitivity and it uses its privacy budget only when a query is above the desired threshold. For our use case, the maximum flippancy can double at most $\log T$ times. Consider first the weaker guarantee of event-level privacy. The maximum flippancy has sensitivity at most $1$ for all event-neighboring streams and the sparse vector algorithm can be used as described. However, for item-neighboring streams, the maximum flippancy can change by $\Omega(T)$ for neighboring streams, whereas the sparse vector algorithm provides good utility only for queries of low sensitivity. The key insight is to ask another type of query: the number of items in the stream with flippancy above the current flippancy bound. This query has sensitivity at most one for item-neighboring streams, as desired. However, it is not clear how to use such queries to estimate a good upper bound on the maximum flippancy of the stream. This is remedied by observing that Algorithm 1, invoked with a flippancy bound $w$, has the same error guarantee even if at most $\sqrt{w}$ items in the stream have flippancy higher than $w$–i.e., an exact upper bound on the maximum flippancy is not needed to design an accurate mechanism. Items that violate the flippancy bound are safely ignored by Algorithm 1 and do not contribute to the distinct elements count. With these ideas, we obtain a mechanism with a $\sqrt{w}$ dependence in the error for the stronger item-level DP guarantee.
@@
Our lower bounds use the embedding technique recently introduced by Jain et al. [43] to obtain strong separations between the batch and continual release models of differential privacy. The approach of Jain et al. embeds multiple separate instances of an appropriately chosen base problem on the same sensitive dataset in the batch model into a single instance of a continual release problem. In this way, the continual release mechanism can be used to solve multiple instances of the base problem in the batch model. The hardness results in the continual release model follow from lower bounds for the batch model.
A key idea in our event-level lower bound is a connection between the inner product of two binary vectors and the number of distinct elements in the union of those indices where the vector bits equal $1$. Estimates of distinct element counts can thus be used to to estimate answers of inner product queries to a sensitive dataset of binary bits. Lower bounds on the accuracy of private inner product queries have been previously established in the batch model through the reconstruction attack of Dinur and Nissim [21]. This connection was used by Mir et al. [49] to provide lower bounds for pan-private algorithms for $\mathsf{CountDistinct}$, however, continual release and pan-privacy are orthogonal notions and their results don’t imply any lower bounds in our setting. We crucially use deletions to embed multiple instances of inner product queries into a stream: once a query is embedded and the desired estimate is received, the elements inserted to answer that query can be entirely deleted from the stream to obtain a “clean slate” for the next query. Thus we obtain a lower bound of $T^{1/4}$ on the error of event-level private mechanisms for $\mathsf{CountDistinct}$ in turnstile streams.
We obtain our stronger item-level lower bounds (for pure and approximate DP) by embedding multiple instances of a $1$-way marginal query. We then apply lower bounds of Hardt and Talwar [39] and Bun et al. [12] for releasing all 1-way marginals in the batch model in conjunction with our reduction. The 1-way marginals of a dataset $y\in\{0,1\}^{n\times d}$, consisting of of $n$ records and $d$ attributes, are the averages of all $d$ attributes of $y$. Deletions in the stream are once again crucially used to embed a marginal query for one attribute and then clean the slate for the next attribute. Changing one record/row in the dataset $y$ translates to $d$ changes of an item in the constructed stream, and thus this reduction is particularly tailored to item-level lower bounds.
@@ 1.3 Related work
The study of differentially privacy in the model of continual release was initiated by two concurrent works [26, 13]. They proposed the binary tree mechanism for computing sums of binary bits. The binary tree mechanism has found numerous applications both in the continual release setting and elsewhere, demonstrating the versatility of this mechanism. Under continual release, it has been extended to work for sums of real values [51], weighted sums [9], graph statistics [54, 32], and most relevantly, counting distinct elements [9, 30, 36]. It has also been employed in the context of private online learning [44, 56, 1] and answering range queries [26, 27, 29].
@@
Prior to our work, the distinct elements problem with continual release has been exclusively studied in the insertions-only model. Bolot et al. [9] were the first to study this problem and showed a $O(\log^{1.5}T)$-accurate item-level-DP mechanism. Ghazi et al. [36] consider additionally the more challenging sliding-window model and show nearly-matching upper and lower bounds for this setting, parameterized by the window size, for item-level and event-level DP. Epasto et al. [30] study the more general $\ell_{p}$-frequency estimation problem with a focus on space efficiency. For distinct elements, i.e., $p=0$, their mechanism provides an estimate with $1+\eta$ multiplicative error and $O(\log^{2}T)$ additive error, using space $\mathrm{poly}(\log T/\eta)$. They also extend their results to the sliding-window model. Two of the works [9, 30] reduce the distinct elements problem to the bit summation primitive, which allows them to use the binary tree mechanism. Since the streams are restricted to be insertion-only, the bit summation primitives they consider have low constant sensitivity. The same primitives have sensitivity $\Omega(T)$ for turnstile streams, and thus this approach cannot be extended to our setting. Ghazi et al. [36] observe that for fixed and sliding windows, the distinct elements problem can be reduced to range queries. For the special case when the window is the entire stream, their reduction is to the summation problem.
@@
Another line of work investigated private sketches for distinct elements, motivated by the popularity of sketching algorithms for the streaming setting. Mergeable sketches for counting distinct elements have received particular attention [55, 16, 50, 40], since they allow multiple parties to estimate the joint number of distinct elements by merging their private sketches. While, these sketches can be combined with the binary tree mechanism to obtain private mechanisms for distinct elements, the utility deteriorates when many $(\log T)$ sketches are merged. In fact, Desfonatines et al. [19] show that achieving both privacy and high accuracy is impossible when many sketches for counting distinct elements are merged. Other private sketches have been studied [53, 20, 57] for the streaming batch setting (without continual release). The distinct elements problem has also been studied in a distributed setting [15, 35], and under pan-privacy [49]. In particular, our lower bound for event-level privacy uses ideas from the lower bound of Mir et al. [49], as described in Section 1.2
@@
The distinct elements problem has been extensively studied in the non-private streaming setting, where the goal is to achieve low space complexity [33, 3, 17, 37, 38, 6, 5, 22, 42, 58, 31, 7, 34, 10, 45]. Blocki et al. [8] show a black-box transformation for every streaming algorithm with tunable accuracy guarantees into a DP algorithm with similar accuracy, for low sensitivity functions. Their transformation does not obviously extend to the continual release setting, and moreover $\mathsf{CountDistinct}$ for turnstile streams does not satisfy the low-sensitivity property.
@@
The first lower bound in the continual release model of differential privacy was an $\Omega(\log T)$ bound on the accuracy of mechanisms for bit summation, shown by Dwork et al [26]. Jain et al. [43] gave the first polynomial separation in terms of error between the continual release model and the batch model under differential privacy. Our lower bounds also show such a separation. The lower bounds of Jain et al. [43] were for the problems of outputting the value and index of the attribute with the highest sum, amongst $d$ attributes of a dataset. Our lower bounds are inspired by their sequential embedding technique to reduce multiple instances of a batch problem to a problem in the continual release model. Similar to them, we also reduce from the 1-way marginals problem to obtain our item-level lower bound. However, our event-level lower bound involves reducing from a different problem, and our reductions use the specific structure of the distinct elements problem for turnstile streams.
@@ 1.4 Broader impact, limitations, and open questions
We study the achievable error of DP mechanisms for counting distinct elements under continual observation in streams with insertions and deletions. We show that it is characterized by the maximum flippancy of the stream. Our work is motivated by societal concerns, but focused on fundamental theoretical limits. It contributes to the broader agenda of obtaining privacy-preserving algorithms for data analysis. We discuss natural directions for future research and some limitations of our work.
Tight Bounds: We have found the best possible error in some settings, but there are parameter regimes where there is a gap between our upper and lower bounds. What is the right error bound for event-level privacy for streams with maximum flippancy $w$ between $\sqrt{T}$ and $T^{2/3}$? Our results yield a lower bound of $T^{1/4}$ and an upper bound of roughly $\sqrt{w}$.
Bounded Memory: We did not consider any memory restrictions (since we focused on purely additive error). It is not clear how to apply the sketching techniques of Epasto et al. [30] to the turnstile setting. It would be interesting to come up with accurate, private, and low-memory mechanisms for counting distinct elements in turnstile streams.

### v3

@@ 1.2 Our techniques
Before describing our algorithmic ideas, we explain the main obstacle to using the techniques previously developed for insertion-only streams [9, 29] in the turnstile setting. Bolot et al. [9] and Epasto et al. [29] used a reduction from $\mathsf{CountDistinct}$ to the summation problem. A mechanism for the summation problem outputs, at every time step $t\in[T]$, the sum of the first $t$ elements of the stream. Dwork et al. [27] and Chan et al. [13] designed the binary-tree mechanism to obtain a $O(\operatorname{polylog}T)$-accurate mechanism for summation. Given an input stream $x$ of length $T$ (to the $\mathsf{CountDistinct}$ problem), define a corresponding summation stream $s_{x}\in\{-1,0,1\}^{T}$. At time step $t\in[T]$, the entry $s_{x}[t]$ equals the difference in the count of distinct elements between time steps $t-1$ and $t$, i.e., $s_{x}[t]=\mathsf{CountDistinct}(x)[t]-\mathsf{CountDistinct}(x)[t-1]$. Then $\mathsf{CountDistinct}(x)[t]$ is precisely the sum of the first $t$ elements of $s_{x}$. In the insertion-only model, changing one entry of $x$ changes at most $2$ entries of $s_{x}$, and thus, by group privacy, the binary-tree mechanism has $O(\operatorname{polylog}T)$ additive error for $\mathsf{CountDistinct}$. For turnstile streams, even under the weaker notion of event-level privacy, a change in the stream $x$ can cause $\Omega(T)$ changes to $s_{x}$. To see this, consider the stream consisting of alternating insertions ($+u$) and deletions ($-u$) of a single element $u\in\mathcal{U}$, and its event-neighboring stream where the first occurrence of $+u$ is replaced with $\perp$. This example illustrates that one of the difficulties of the $\mathsf{CountDistinct}$ problem for turnstile streams lies with items that switch from being present to absent multiple times over the course of the stream, that is, have high flippancy. We present a private mechanism that outputs estimates of the count of distinct elements in a turnstile stream with optimal accuracy in terms of maximum flippancy.
@@
Our first key idea allows us to obtain a mechanism, Algorithm 1, that is given as input a flippancy upper bound $w$. For a stream $x$ whose maximum flippancy is bounded by $w$, changing to an item-neighbor of $x$ (whose maximum flippancy is also bounded by $w$) causes at most $2w$ changes to the corresponding summation stream $s_{x}$. This observation, combined with a group privacy argument, gives a mechanism with error $O(w\cdot\operatorname{polylog}\ T)$ directly from the accuracy guarantee of the binary-tree mechanism for summation. Previous works in the insertion-only model [9, 29] used precisely this approach for the special case $w=1$. To obtain the better $\sqrt{w}$ dependence on $w$ in our upper bound, we “open up” the analysis of the binary-tree mechanism. By examining the information stored in each node of the binary tree for the summation stream, we show that changing the occurrences of one item in a stream $x$ with maximum flippancy at most $w$ can change the values of at most $w$ nodes in each level of the binary tree. The $\sqrt{w}$ dependence in the error then follows from the privacy guarantees of the Gaussian mechanism (used inside the binary-tree mechanism) for approximate differential privacy. This type of noise reduction makes crucial use of the binary tree approach: there are optimized noise addition schemes for prefix sums that improve quantitatively over the binary-tree mechanism (see, e.g., [18, 41]), but it is unclear if they allow the same noise reduction. While our mechanism is only accurate for streams with maximum flippancy at most $w$, it is private even for streams that violate this condition. To achieve this, our mechanism ignores stream elements after their flippancy exceeds $w$.
The second key idea allows our algorithms to adapt automatically to the maximum flippancy $w_{x}$ of the input, without the need for an a-priori bound $w$. We design a private mechanism, Algorithm 3, that approximately keeps track of the maximum flippancy of the prefix of the stream seen so far and invokes our first mechanism (Algorithm 1) with the current estimated maximum flippancy $\hat{w}$ as an input. Our main innovation lies in the careful application of the sparse vector algorithm [24] to track the maximum flippancy of the stream. We cannot do this directly, since the sparse vector algorithm achieves good utility only for queries of low sensitivity, and maximum flippancy has global sensitivity $\Omega(T)$ under both event-level and item-level changes.
@@
Instead, we track a low sensitivity proxy that indirectly monitors the maximum flippancy $w_{x}$: given the current estimate $\hat{w}$ of the flippancy, we use the sparse vector algorithm to continuously query the number of items in the stream with flippancy above $\hat{w}$. This query has sensitivity one for item-level neighbors, as desired, but it is not a priori clear how to use it to upper bound the maximum flippancy of the stream. This is remedied by observing that Algorithm 1, invoked with a flippancy bound $\hat{w}$, has the same error (and privacy) guarantee even if at most $\sqrt{\hat{w}}$ items in the stream have flippancy higher than $\hat{w}$. That is, an exact upper bound on the maximum flippancy is not needed to design an accurate mechanism. Items that violate the flippancy bound are safely ignored by Algorithm 1 and do not contribute to the distinct elements count.
When the number of high-flippancy items gets large, we adjust the estimate $\hat{w}$ and invoke a new instantiation of Algorithm 1. By doubling $\hat{w}$ each time this happens, we ensure that it remains at most twice the actual maximum flippancy $w_{x}$, and that we need only invoke $\log T$ different copies of Algorithm 1 and the sparse vector algorithm [^1] All $\log$ expressions in this article are base $2$.. With these ideas, we obtain an item-level private mechanism that, for all streams $x$, has error that scales with $\sqrt{w_{x}}$.
@@
Our lower bounds use the embedding technique introduced by Jain et al. [44] to obtain strong separations between the batch and continual release models of differential privacy. The approach of Jain et al. embeds multiple separate instances of an appropriately chosen base problem on the same sensitive dataset in the batch model into a single instance of a continual release problem. Then, the continual release mechanism can be used to solve multiple instances of the base problem in the batch model. The hardness results in the continual release model follow from lower bounds for the batch model.
A key idea in our event-level lower bound is a connection between the inner product of two binary vectors and the count of distinct elements in the union of those indices where the vector bits equal $1$. Estimates of distinct elements counts can thus be used to estimate inner products on a sensitive dataset of binary bits. Lower bounds on the accuracy of private algorithms for estimating inner product queries have been previously established in the batch model through the reconstruction attack of Dinur and Nissim [21]. This connection was used by Mir et al. [50] to provide lower bounds for pan-private algorithms for counting distinct elements. However, continual release and pan-privacy are orthogonal notions, and their results don’t imply any lower bounds in our setting. We crucially use deletions to embed multiple instances of inner product queries into a stream: once a query is embedded and the desired estimate is received, the elements inserted to answer that query can be entirely deleted from the stream to obtain a “clean slate” for the next query. We obtain a lower bound of $T^{1/4}$ on the error of event-level private mechanisms for $\mathsf{CountDistinct}$ in turnstile streams.
We obtain our stronger item-level lower bounds (for pure and approximate differential privacy) by embedding multiple instances of a $1$-way marginal query. We then apply lower bounds of Hardt and Talwar [38] and Bun et al. [12] for releasing all 1-way marginals in the batch model in conjunction with our reduction. The 1-way marginals of a dataset $y\in\{0,1\}^{n\times d}$, consisting of $n$ records and $d$ attributes, are the averages of all $d$ attributes of $y$. Deletions in the stream are once again crucially used to embed a marginal query for one attribute and then clean the slate for the next attribute. Changing one record/row in the dataset $y$ translates to $d$ changes of an item in the constructed stream, and thus this reduction is particularly tailored to item-level lower bounds.
@@ 1.3 Related work
The study of differential privacy under continual release was initiated by two concurrent works [25, 13]. They proposed the binary-tree mechanism for computing sums of binary bits. The versatility of this mechanism is demonstrated by numerous applications it has found in the continual release setting and elsewhere. Under continual release, it has been extended to work for sums of real values [52], weighted sums [9], graph statistics [55, 31], and most relevantly, counting distinct elements [9, 29, 35]. It has also been employed for private online learning [45, 57, 1] and for answering range queries [25, 26, 28].
@@
Prior to our work, the $\mathsf{CountDistinct}$ problem with continual release was studied exclusively in the insertions-only model. Bolot et al. [9] were the first to study this problem and showed a $O(\log^{1.5}T)$-accurate item-level-DP mechanism. Ghazi et al. [35] considered the more challenging sliding-window model and showed nearly-matching upper and lower bounds for this setting, parameterized by the window size, for item-level and event-level differential privacy. Epasto et al. [29] studied the more general $\ell_{p}$-frequency estimation problem with a focus on space efficiency. For distinct elements, i.e., $p=0$, their mechanism provides an estimate with $1+\eta$ multiplicative error and $O(\log^{2}T)$ additive error, using space $\mathrm{poly}(\log T/\eta)$. They also extended their results to the sliding-window model. Two of the works [9, 29] reduced the $\mathsf{CountDistinct}$ problem to the bit summation primitive, which allowed them to use the binary-tree mechanism. Since the streams are restricted to be insertion-only, the bit summation primitives they considered have low constant sensitivity. The same primitives have sensitivity $\Omega(T)$ for turnstile streams, and thus this approach cannot be directly extended to our setting. Ghazi et al. [35] observed that for fixed and sliding windows, the distinct elements problem can be reduced to range queries. For the special case when the window is the entire stream, their reduction is to the summation problem.
In concurrent work, Henzinger et al. [40] studied $\mathsf{CountDistinct}$ with insertions and deletions in the likes model (see “Variants of the model” in Section 1.1). Our model is more general, since it allows for multiple consecutive insertions and deletions of the same item. Our upper bound and our item-level privacy lower bound can be extended to the likes model. In contrast, our event-level private lower bound provably does not apply to that model: in the likes model, for event-level privacy, there is a simple reduction to the bit summation problem in the continual release model such that the resulting algorithm incurs only a polylogarithmic in $T$ error, whereas we show that in our model, any event-level private algorithm incurs a polynomial in $T$ error.
Henzinger et al. [40] showed error bounds for item-level privacy in the likes model that are parameterized by the total number of updates $K$ in the stream. The parameter $K$ is related to our concept of flippancy: in the likes model, $K$ equals the sum of all items’ flippancies and, in general, is at least that sum. Henzinger et al. [40] give an $(\varepsilon,0)$-DP algorithm with error $\tilde{O}(\sqrt{K}\log T)$ and show a nearly matching lower bound on the error for $(\varepsilon,0)$-DP algorithms using a packing argument. This lower bound applies to our model as well. It is incomparable to our lower bounds, since it scales differently and depends on a different parameter. In our model, their algorithm can be analyzed to give error bounds in terms of the sum $K^{\prime}$ of the flippancies of the items and incurs error $\tilde{O}(\sqrt{K^{\prime}}\log T)$; however, it is unclear if their algorithm can be analyzed in our model to give bounds in terms of the (potentially smaller) maximum flippancy.
@@
Another line of work investigated private sketches for distinct elements, motivated by the popularity of sketching algorithms for the streaming setting. Mergeable sketches for counting distinct elements have received particular attention [56, 15, 51, 39], since they allow multiple parties to estimate the joint count of distinct elements by merging their private sketches. While these sketches can be combined with the binary-tree mechanism to obtain private mechanisms for $\mathsf{CountDistinct}$, the utility deteriorates when many $(\log T)$ sketches are merged. In fact, Desfontaines et al. [19] showed that achieving both privacy and high accuracy is impossible when many sketches for counting distinct elements are merged. Other private sketches have been studied [54, 20, 58] for the streaming batch setting (without continual release). The distinct elements problem has also been studied in a distributed setting [14, 34] and under pan-privacy [50]. In particular, our lower bound for event-level privacy uses ideas from the lower bound of Mir et al. [50], as described in Section 1.2
@@
The $\mathsf{CountDistinct}$ problem has been extensively studied in the non-private streaming setting, where the goal is to achieve low space complexity [32, 3, 16, 36, 37, 6, 5, 22, 43, 59, 30, 7, 33, 10, 46]. Blocki et al. [8] showed a black-box transformation for every streaming algorithm with tunable accuracy guarantees into a DP algorithm with similar accuracy, for low sensitivity functions. Their transformation does not obviously extend to the continual release setting. Moreover $\mathsf{CountDistinct}$ has high sensitivity for turnstile streams.
@@
The first lower bound in the continual release model of differential privacy was an $\Omega(\log T)$ bound on the accuracy of mechanisms for bit summation, shown by Dwork et al. [25]. Jain et al. [44] gave the first polynomial separation in terms of error between the continual release model and the batch model under differential privacy. Our lower bounds also show such a separation. The lower bounds of Jain et al. [44] were for the problems of outputting the value and index of the attribute with the highest sum, amongst $d$ attributes of a dataset. Our lower bounds are inspired by their sequential embedding technique to reduce multiple instances of a batch problem to a problem in the continual release model. Similar to them, we also reduce from the 1-way marginals problem to obtain our item-level lower bound. However, our event-level lower bound involves reducing from a different problem, and our reductions use the specific structure of $\mathsf{CountDistinct}$ for turnstile streams.
@@ 1.4 Broader impact, limitations, and open questions
We study the achievable error of DP mechanisms for $\mathsf{CountDistinct}$ under continual observation in streams with insertions and deletions. We show that it is characterized by the maximum flippancy of the stream. Our work is motivated by societal concerns, but focused on fundamental theoretical limits. It contributes to the broader agenda of obtaining privacy-preserving algorithms for data analysis. We discuss natural directions for future research and some limitations of our work.
Tight bounds: We found the best achievable error in some settings, but our upper and lower bounds do not match in some parameter regimes. What is the right error bound for event-level privacy for streams $x$ with maximum flippancy $w_{x}$ between $\sqrt{T}$ and $T^{2/3}$? Our results yield a lower bound of $T^{1/4}$ and an upper bound of roughly $\sqrt{w_{x}}$.
Bounded memory: We did not consider any memory restrictions. Prior to our work, no other work addressed $\mathsf{CountDistinct}$ with deletions under continual release—with or without space constraints. We consider only the privacy constraint since it is more fundamental—it cannot be avoided by buying more memory—and the best algorithms with unbounded memory provide a benchmark by which to evaluate space-constrained approaches.
Space complexity is certainly a natural topic for future work. While it is not clear how to apply the sketching techniques of Epasto et al. [29] to the turnstile setting, it would be interesting to come up with accurate, private, and low-memory mechanisms for counting distinct elements in turnstile streams. Such algorithms would necessarily mix multiplicative and additive error guarantees (due to space and privacy constraints, respectively).